15. Find the standard deviation of the penny date data, and determine the z-score for the dates of 2012 and 2009. What is the significance of the sign of the z-score of each?
Charles has a jar in which he places any pennies that he may obtain during his daily activities. His sister, Oluoma, takes a handful of pennies off the top and records the dates on the pennies: 2013, 23014s 2oos, 212, 2, 2013,2012,2013,.2011.2011, 2010,2003, 212. 2013, 2012, 2010.
Answer #1

found these stats using MINITAB

Variable Mean StDev
date 2011.4 1.50

So, Standard deviation is 1.50

So, z score of 2012 = (2012-2011.4)/1.50 = 0.4 (+) sign refers this data is higher than average

and z score of 2009 = (2009-2011.4)/1.5 = -1.6 (-) sign refers this data is lower than average